(1-6) The following table presents the number of reports of graffiti in each of New York’s five boroughs over a one-year period. These reports were classified as being open, closed, or pending.

Borough

Open Report

Closed Report

Pending Report

Total

Bronx

1121

1622

80

2823

Brooklyn

1170

2706

48

3924

Manhattan

744

3380

25

4149

Queens

1353

2043

25

3421

Staten Island

83

118

0

201

Total

4471

9869

178

14518

Compute the following probabilities:

1. Find the probability a random report is open and comes from Brooklyn.

2. Find the probability a random report comes from Manhattan.

3. Find the probability a random report is not from Queens.

4.Given a report is from Staten Island, what is the probability it is closed.

5. Find the probability a random report comes from Brooklyn or is open.

6. Find the probability a report is from Manhattan, given it is pending.
Answer #1

1)

The required probability = 1170/14518 = 0.0806

2)

The required probability = 4149/14518 = 0.2858

3)

The required probability = 1 - (3421/14518) = 0.7644

4)

The required probability = 118/201 = 0.5871

5)

The required probability = P(Brooklyn) + P(Open) - P(Brooklyn and Open)

= (3924 + 4471 - 1170)/14518 = 0.4977

6)

The required probability = 25/178 = 0.1404